OpenStudy (anonymous):

Find the holes of f(x) = (3x^2+2x-1)/(x^2-1)

OpenStudy (anonymous):

Also, may you find the range?

OpenStudy (calculusfunctions):

Step 1: Factor the numerator and the denominator of the rational function, completely. Step 2: Divide factors common to both numerator and denominator. Step 3: Any factors that are completely removed from the denominator, the restriction from these factors, are the x-coordinates of the points where the removable discontinuities (holes) are. Calling them removable discontinuities is much more elegant then calling them holes, don't you think? \[f(x) = \frac{ 3x ^{2} + 2x - 1}{x ^{2} - 1}\] x ≠ -1, 1 Step 1: \[f(x) = \frac{ (3x - 1)(x + 1) }{(x - 1)(x + 1) }\] Step 2: \[f(x) = \frac{ 3x - 1 }{x - 1 }\] Step 3: The factor of (x + 1) was removed from the denominator. Thus there is a removable discontinuity (hole) at x = -1. To find the y-coordinate of the point of removable discontinuity, evaluate the simplified fraction for x = -1. Thus \[f(x) = \frac{ 3x - 1 }{ x - 1 }\] f(-1) = 2 ∴ the removable discontinuity (hole) occurs at the point (-1, 2) You review this while I explain range.

OpenStudy (anonymous):

how did you get 2?

OpenStudy (calculusfunctions):

\[f(x) = \frac{ 3x - 1 }{ x - 1 }\] \[f(-1) = \frac{ 3(-1)-1 }{ -1-1 }\] \[f(-1) = \frac{ -3-1 }{ -1- 1 }\] \[f(-1) = \frac{ -4 }{ -2 }\] f(-1) = 2

OpenStudy (anonymous):

Wait what??? I thought the zeroes were supposed to make the equation equal zero?

OpenStudy (calculusfunctions):

You didn't ask for zeros, you asked where the removable discontinuities are. Zeros are the x-intercepts.

OpenStudy (calculusfunctions):

The x-intercepts occur when f(x) = 0. A rational function can only equal zero if the numerator of the simplified fraction equals zero. Thus in the case of the above the zero(s) (x-int.) is at x = 1/3.

OpenStudy (anonymous):

Wish me luck on the test tomorrow

OpenStudy (calculusfunctions):

I was typing the explanation to find the range and the computer rebooted. I have to start over. Give me a few minutes.

OpenStudy (calculusfunctions):

To find the range algebraically, write the simplified equation in terms of x (solve the equation for x). \[f(x) = \frac{ 3x -1 }{ x -1 }\] since y = f(x) \[y = \frac{ 3x -1 }{ x -1 }\] Multiply both sides of the equation by (x - 1) to obtain y(x - 1) = 3x - 1 xy - y = 3x - 1 xy - 3x = y - 1 Since we need to solve for x, factor x from the left side of the equation. x(y - 3) = y - 1 Divide both sides by (y - 3). to obtain \[x = \frac{ y -1 }{ y -3 }\] The restriction here is y ≠ 3. ∴ the range of f(x) is {y ∈ ℝ | y ≠ 3}. Please let me know if this helps since your test is tomorrow, and GOOD LUCK!!